OpenStudy (anonymous):

Simplify the complex rational expression. 1/x + 6 / 5/x^2 - 36

OpenStudy (anonymous):

@satellite73 could you help me please?

OpenStudy (anonymous):

@Hero could you help me please?

OpenStudy (anonymous):

@agent0smith could you help me with this one please?

OpenStudy (agent0smith):

This is hard to read. Use parentheses. \[\huge \frac{ 1 }{x } + \frac{ 6 }{ \frac{ 5 }{ x^2 - 36 }}\] no idea.

OpenStudy (agent0smith):

If that's what it is, then simplify the right fraction... since you're dividing by a fraction, you can multiply by the reciprocal.

OpenStudy (anonymous):



OpenStudy (anonymous):

thats how its written

OpenStudy (anonymous):

the choices are A. 5/ x - 6 B. x - 6 / 5 C. x + 6 / 5 D . x - 6

OpenStudy (agent0smith):

Okay. This is why you need to use parentheses, or draw it, or something :P because this "1/x + 6 / 5/x^2 - 36" looks nothing like that. Since you're diving by a fraction, you can multiply by the reciprocal: \[\huge \frac{ \frac{ 1 }{x + 6 } }{ \frac{ 5 }{ x^2 - 36 } } = \] \[\huge \frac{ 1 }{(x + 6) } \times \frac{ (x^2 - 36) }{ 5 } = \]

OpenStudy (anonymous):

sorry about that could you break it down more

OpenStudy (agent0smith):

Factor x^2-36, it's a difference of two squares.